OpenStudy (anonymous):

If 4.80g of CuNO3 is dissolved in water to make 0.270M. what is the volume of the solution?

OpenStudy (zpupster):

molar mass of CuNO3 =63.5 +14+3*16=125.5g/mole .27 M = .27 moles/L (.27 moles/L)(125.5g/mole) = X grams/L then take 4.8g/ (X grams/L) = this answer will be in Liters --watch the units, it will help you solve chem problems.

OpenStudy (anonymous):

That didn't work

OpenStudy (anonymous):

I fixed it. It's suppose to be .270